OpenStudy (anonymous):

Mrs. Winston sells chocolate fudge for $7.50 per pound and peanut butter fudge for $7.00 per pound. The total number of pounds sold on saturday was 146 and the total amount of money collected was $1065. How many pounds of each type of fudge were sold?

OpenStudy (kropot72):

Let C be the number of pounds of chocolate fudge that was sold, and let P be the number of pounds of peanut butter fudge that was sold. Now we can write the following equations: C + P = 146 ................(1) 7.5C + 7P = 1065 .......(2) From equation (1) we can find P in terms of C by rearranging as follows: P = 146 - C ...............(3) Next we need to plug the value for P in (3) into equation (2) and solve the result to find the value of C.

OpenStudy (anonymous):

 this will work for problems similar to this